OpenStudy (anonymous):

8^(x-2)=5^x medal/follow

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

take the log of both sides and use some log laws \[\log(8^{x -2}) = \log(5^x)\] applying the log laws for powers \[(x - 2)\log(8) = x \log(5)\] so distributing and collecting like terms \[x(\log(8) - \log(5)) = 2\log(8)\] so \[x = \frac{2\log(8)}{\log(8) - \log(5)}\] hope it helps

OpenStudy (anonymous):

thanks so much